OpenStudy (coconutjj):

Integrate(2x^-5). How do you do this question?

OpenStudy (solomonzelman):

is it \(\large\color{black}{\displaystyle\int\limits_{~}^{~}2x^{-5}dx}\) ?

OpenStudy (coconutjj):

yes

OpenStudy (solomonzelman):

then you can take the constant out, like this: 2\(\large\color{black}{\displaystyle\int\limits_{~}^{~}x^{-5}dx}\) and apply the power rule.

OpenStudy (solomonzelman):

\(\large\color{black}{2\displaystyle\int\limits_{~}^{~}x^{-5}dx=2\times \frac{\LARGE x^{-5\color{red}{+1}} }{\LARGE \color{red}{-5+1}}+C}\)

OpenStudy (solomonzelman):

what do you get as your answer?

OpenStudy (coconutjj):

(2(x^-4)/(-4))

OpenStudy (solomonzelman):

yes, right: \(\large\color{black}{ 2 \times \frac{\LARGE x^{-4} }{\LARGE -4} +C }\) but you can simplify this a bit more

OpenStudy (coconutjj):

-(x^-4/2)

OpenStudy (solomonzelman):

yup, \(\large\color{black}{ \frac{\LARGE x^{-4} }{\LARGE -2} +C }\)

OpenStudy (coconutjj):

thank you..

OpenStudy (solomonzelman):

or, \(\large\color{black}{ -\frac{\LARGE x^{-4} }{\LARGE 2} +C }\) same thing

OpenStudy (solomonzelman):

what does your teacher require to to farther do, does (s)he want you to re-write using a positive exponent?

OpenStudy (solomonzelman):

or would this result satisfy your teacher?

OpenStudy (coconutjj):

Yes this result works

OpenStudy (solomonzelman):

okay, but make sure to include the +C :)

OpenStudy (solomonzelman):

You\(\normalsize\color{blue}{ \rm ~welcome! }\)